AC309 Unit 5: Application Assignment – Investments

Millie is a single taxpayer and her 2018 taxable income is $181,205, calculated as follows:

Description

Amount

Employee wages

133,000

Profit from business (Schedule C)

15,000

Ordinary income from partnership (Schedule E)

22,000

Interest income

4,000

Ordinary dividends (total dividends = 9,000)

2,000

Qualified dividends (total dividends = 9,000)

7,000

Short-term capital gain

8,200

Long-term capital gain

6,900

Total Income – Form 1040, Line 6

198,100

Deduction for ½ of self-employment tax

(495)

Adjusted Gross Income – Form 1040, Line 7

197,605

Standard deduction – Form 1040, Line 8

(12,000)

Qualified business income deduction – Form 1040, Line 9

(4,400)

Taxable Income – Form 1040, Line 10

181,205

Ordinary Income Bracket

Tax Calculation

$0 - $9,525

10% of ordinary income

$9,526 - $38,700

$952.50 plus 12% of ordinary income over $9,525

$38,701 - $82,500

$4,453.50 plus 22% of ordinary income over $38,700

$82,501 - $157,500

$14,089.50 plus 24% of ordinary income over $82,500

$157,501 - $200,000

$32,089.50 plus 32% of ordinary income over $157,500

$200,001 - $500,000

$45,689.50 plus 35% of ordinary income over $200,000

$500,001 or more

$150,689.50 plus 37% of ordinary income over $500,000

Instructions

  1. Separate Millie’s 2018 taxable income of $181,205 into its ordinary income and preferential income components. Hint – your ordinary income and preferential income amounts should total Millie’s taxable income of $181,205.

Taxable Income

Ordinary Income

Preferential Income

181,205

  1. Using the single taxpayer ordinary tax rate brackets on the previous page, compute Millie’s 2018 income tax liability she should report on her 2018 Form 1040, Line 11. Do NOT calculate any self-employment tax, additional Medicare tax or net investment income tax that Millie might also be liable for (these additional taxes would be reported on Schedule 4 and Form 1040, Line 14).

Description

Amount

Tax liability – ordinary income

Tax liability – preferential income

2018 income tax liability (Form 1040, Line 11)
Answer #1

Taxable income

Ordinary income

Preferential income

181205

$167305

$13900

Preferential income = qualified dividends + long term capital gain = 7000+6900 = 13900

Ordinary income = total taxable income -13900 = 181205-13900 = $167305

Description

Amount

Tax liability – ordinary income (32089.50+(32%*(167305-157500)))

$35227

Tax liability – preferential income (13900*15%)

2085

2018 income tax liability (Form 1040, Line 11)

$37312

As taxable income is between $38,700-$426,700, the applicable preferential tax rate is 15%